OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (4x - 7) units and the length of side ON is (7 – 3x) units. What value of x will make it a rhombus? 14 2 7 12

OpenStudy (campbell_st):

a rhombus has equal side lengths... so 4x - 7 = 7 - 3x add 3x to moth sides 7x - 7 = 7 you can finish it from here

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

so it's 7 @campbell_st

OpenStudy (campbell_st):

add 7 to both sides 7x = 14

OpenStudy (anonymous):

2 :D

OpenStudy (campbell_st):

thats it